OpenStudy (anonymous):

What is the center and radius of the circle with the equation x2 + y2 + 8x – 12y + 36 = 0 ? Center: (5, 6); Radius = 4 Center: (–4, 6); Radius = 5 Center: (5, 6); Radius = 5 Center: (–4, 6); Radius = 4

OpenStudy (campbell_st):

complete the square for x and y x^2 + 8x + y^2 - 12y = -36 x^2 + 8x + 16 + y^2 - 12y + 36 = -36 + 36 + 16 factorising (x +4)^2 +(y -6)^2 = 4^2 the centre is (-4, 6) radius 4

OpenStudy (campbell_st):

the general form of a circle is (x - h)^2 + (y - k)^2 = r^2 centre (h, k) radius = r

OpenStudy (anonymous):

Thanks (: